# VideoPASTA [Uncaptioned image]: 7K Preference Pairs That Matter for Video-LLM Alignment

### v1

@@ A.3.1 Dataset Statistics
Starting with 3000 videos from ActivityNet [64], we systematically generate preference pairs through structured adversarial sampling. For each video $V$, we generate 10 queries $Q$ targeting different aspects of video understanding. Each query $q\in Q$ is paired with three targeted adversarial responses $r_{\text{spatial}}$, $r_{\text{temporal}}$, and $r_{\text{crossframe}}$, representing spatial, temporal, and cross-frame failure modes, respectively. Theoretically, this setup yields:
@@
However, to ensure dataset quality, we employ rigorous filtering using Qwen2.5-32B [61] verification using the prompt template given in Figure 13. Each preference pair must satisfy three criteria:
@@
    The preferred response should accurately reflect the video content relative to the query.
@@
This verification process retains approximately 7.8% of the potential pairs:
$N_{\text{final}}=N_{\text{potential}}\times r_{\text{retention}}\approx 90\,000\times 0.078\approx 7,020,$ (6)
where $r_{\text{retention}}$ is the retention rate after quality filtering.
This filtered dataset provides a balanced representation across failure modes while maintaining high standards for preference pair quality. The strict filtering ensures that each adversarial example presents a genuine challenge for video-language alignment rather than simple errors or rephrasing.
A.3.2 Dataset Samples
Figure 8: Examples from our preference dataset demonstrating diverse failure modes targeted by VideoPASTA. Each row shows a video sequence with its corresponding query and preferred and adversarial responses. The adversarial samples cover spatial misalignment (counting objects in scenes), temporal incoherence (order of actions in cooking/preparation), and cross-frame disconnection (scene transitions and contextual shifts). Adversarial responses deliberately introduce specific misalignments by either negating obvious visual elements, confusing sequential ordering, or collapsing distinct temporal phases into simultaneous events, while preferred responses maintain accurate spatial-temporal alignment with the video content.
Figure 8 illustrates key examples from our preference dataset that demonstrate how VideoPASTA targets specific failure modes in video-language understanding. These examples were carefully curated to challenge different aspects of video comprehension while maintaining clear distinctions between preferred and dispreferred responses.
@@ A.3.2 Dataset Samples
The boat counting example demonstrates our approach to spatial reasoning. While the dispreferred response completely negates the presence of obvious visual elements (“no boats”), the challenge lies not in the simple presence/absence but in the precise spatial relationships (“positioned near the shore, with one slightly further out”). This forces the model to develop fine-grained spatial awareness rather than just object detection capabilities.
@@
Two examples highlight our approach to temporal understanding. The cooking sequence tests precise transitional timing between steps, where the dispreferred response artificially collapses distinct preparation phases into simultaneous actions. Similarly, the equipment preparation example challenges the model’s ability to distinguish between sequential and concurrent actions. These adversarial samples are particularly effective because they present plausible but incorrect temporal relationships.
@@
The scene transition example illustrates how we assess long-range comprehension. The dispreferred response mistakenly interprets superficial visual changes, such as a close-up of a face, as significant narrative shifts, whereas the preferred response accurately identifies meaningful context transitions, like an external threat leading to an internal response. This evaluates the model’s ability to track narrative progression across distant frames.
Each example undergoes thorough validation using Qwen2.5-32B [61] to ensure that dispreferred responses reflect genuine misunderstandings rather than simple errors or rephrasings. This systematic approach to adversarial example generation reinforces robust video-language alignment across multiple dimensions of video understanding.
A.4 Qualitative Examples
@@
First, in the camera advertisement sequence, while Qwen2.5-VL [3] fails to recognize the narrative structure and describes it as “unrelated clips” VideoPASTA successfully captures the purposeful progression from technical camera operation to creative photography. This demonstrates how our cross-frame adversarial sampling helps the model develop a more coherent understanding of extended narratives. Next, the animated sequence with Bugs Bunny showcases VideoPASTA’s enhanced ability to track emotional progression. Instead of merely detecting immediate reactions, our model recognizes the escalation from initial irritation to visible anger and, ultimately, to explosive rage. This improvement stems from our temporal incoherence adversarial sampling, which teaches the model to distinguish between simultaneous and sequential emotional states. The cooking demonstration particularly highlights the benefits of our local spatial alignment strategy. While the baseline model confuses the order of preparation steps, VideoPASTA correctly identifies the precise sequence of cleaning, coating, and frying the chilies. This accuracy in tracking procedural steps is crucial for practical applications like instructional video understanding. The competition example shows how our model can parse complex sequences of physical and emotional reactions, maintaining temporal coherence even in dynamic scenes. The eclipse footage example reveals VideoPASTA’s ability to describe gradual visual transformations accurately, avoiding the baseline’s tendency to oversimplify temporal transitions. Finally, the instruction scene identifying magic demonstrates our model’s capability to establish clear causal relationships between actions and their outcomes, supported by specific visual evidence.
@@ A.5 Prompt Templates
Preference Data Filtering Prompt You have a single video input and a set of four responses for each query: 1. One preferred response that is claimed to be well-aligned with the video content. 2. Three adversarial responses, each intentionally introducing spatial, temporal, or cross-frame errors. The goal is to validate that: The preferred response truly aligns with the query (no unintended contradictions or inaccuracies). Each adversarial response introduces a clear misalignment without merely restating or slightly rephrasing the preferred. For each query and its four responses: 1. Sanity-check the preferred response. Confirm that it accurately reflects the video’s content in relation to the query. If any errors or contradictions are detected, discard them. 2. Examine each adversarial response. Identify whether it deliberately contradicts or distorts the query/video content (e.g., reversed sequence, false spatial claims). If it is too similar to the preferred response or fails to demonstrate a clear misalignment, discard it.
Figure 13: Prompt template for validating one preferred and three adversarial responses to ensure robust preference pairs.
@@
Adversarial Example Evaluation Prompt Task: Evaluate whether the provided adversarial example correctly targets its intended failure mode in video understanding. Query: [Original question asked about the video] Preferred Response: [The correct/preferred response to the query] Adversarial Example: [The adversarial example to be evaluated] Claimed Failure Mode: [One of: “Spatial Misalignment”, “Temporal Incoherence”, or “Cross-Frame Disconnection”] Failure Mode Definitions: Spatial Misalignment: Incorrectly describing spatial relations, object positions, occlusion patterns, depth, or relative positioning within a single frame. Temporal Incoherence: Violating the natural ordering of events, describing sequential actions as simultaneous, merging distinct events, or misordering the sequence of activities shown in the video. Cross-Frame Disconnection: Breaking object persistence across frames, describing the same object as different entities across scenes, failing to maintain character/object consistency, or incorrectly describing changes between distant frames. Evaluation Instructions: 1. Carefully analyze the adversarial example in relation to the preferred response. 2. Determine if the adversarial example genuinely induces the claimed failure mode. 3. Your evaluation should be based solely on the definitions provided above. 4. Provide a binary judgment: “Yes” if the adversarial example correctly targets the claimed failure mode, “No” if it does not. 5. Briefly explain your reasoning (2-3 sentences). Output Format:

### v3

@@ E.1 Dataset Statistics
Starting with 3000 videos from ActivityNet Yu et al. (2019), we systematically generate preference pairs through structured adversarial sampling. For each video $V$, we generate 10 queries $Q$ targeting different aspects of video understanding. Each query $q\in Q$ is paired with three targeted adversarial responses $r_{\text{spatial}}$, $r_{\text{temporal}}$, and $r_{\text{crossframe}}$, representing spatial, temporal, and cross-frame failure modes, respectively. Theoretically, this setup yields:
@@
However, to ensure dataset quality, we employ rigorous filtering using Qwen2.5-32B Yang et al. (2024) verification using the prompt template given in Figure 12. Each preference pair must satisfy three criteria:
@@
    The aligned response should accurately reflect the video content relative to the query.
@@
This verification process retains approximately 7.8% of the potential pairs (on LLaVA-NeXT-Interleave):
$N_{\text{final}}$ $=N_{\text{potential}}\times r_{\text{retention}}$ (4)
[…]
where $r_{\text{retention}}$ is the retention rate after quality filtering. This filtered dataset provides a balanced representation across failure modes while maintaining high standards for preference pair quality. The strict filtering ensures that each adversarial example presents a genuine challenge for video-language alignment rather than simple errors or rephrasing.
E.2 Adversarial Sample Diversity
Figure 7: Adversarial example diversity. Each row shows a video sequence with its corresponding query and aligned and adversarial responses. The adversarial samples cover spatial misalignment (counting objects in scenes), temporal incoherence (order of actions in cooking/preparation), and cross-frame disconnection (scene transitions and contextual shifts). Adversarial responses deliberately introduce specific misalignments by either negating obvious visual elements, confusing sequential ordering, or collapsing distinct temporal phases into simultaneous events, while aligned responses maintain accurate spatial-temporal alignment with the video content.
Figure 7 illustrates the diversity of adversarial examples demonstrating how VideoPASTA targets specific failure modes. These examples were carefully curated to challenge different aspects of video comprehension while maintaining clear distinctions between aligned and adversarial responses.
@@ E.2 Adversarial Sample Diversity
The boat counting example demonstrates our approach to spatial reasoning. While the adversarial response completely negates the presence of obvious visual elements (“no boats”), the challenge lies not in the simple presence/absence but in the precise spatial relationships (“positioned near the shore, with one slightly further out”). This forces the model to develop fine-grained spatial awareness rather than just object detection capabilities.
@@
Two examples highlight our approach to temporal understanding. The cooking sequence tests precise transitional timing between steps, where the adversarial response artificially collapses distinct preparation phases into simultaneous actions. Similarly, the equipment preparation example challenges the model’s ability to distinguish between sequential and concurrent actions. These adversarial samples are particularly effective because they present plausible but incorrect temporal relationships.
@@
The scene transition example illustrates how we assess long-range comprehension. The adversarial response mistakenly interprets superficial visual changes, such as a close-up of a face, as significant narrative shifts, whereas the aligned response accurately identifies meaningful context transitions, like an external threat leading to an internal response. This evaluates the model’s ability to track narrative progression across distant frames.
Each example undergoes thorough validation using Qwen2.5-32B Yang et al. (2024) to ensure that adversarial responses reflect genuine misunderstandings rather than simple errors or rephrasings. This systematic approach to adversarial example generation reinforces robust video-language alignment across multiple dimensions of video understanding.
Appendix F Qualitative Examples
@@
First, in the camera advertisement sequence, while Qwen2.5-VL Bai et al. (2025) fails to recognize the narrative structure and describes it as “unrelated clips” VideoPASTA successfully captures the purposeful progression from technical camera operation to creative photography. This demonstrates how our cross-frame adversarial sampling helps the model develop a more coherent understanding of extended narratives. Next, the animated sequence with Bugs Bunny showcases VideoPASTA’s enhanced ability to track emotional progression. Instead of merely detecting immediate reactions, our model recognizes the escalation from initial irritation to visible anger and, ultimately, to explosive rage. This improvement stems from our temporal incoherence adversarial sampling, which teaches the model to distinguish between simultaneous and sequential emotional states. The cooking demonstration particularly highlights the benefits of our local spatial alignment strategy. While the baseline model confuses the order of preparation steps, VideoPASTA correctly identifies the precise sequence of cleaning, coating, and frying the chilies. This accuracy in tracking procedural steps is crucial for practical applications like instructional video understanding. The competition example shows how our model can parse complex sequences of physical and emotional reactions, maintaining temporal coherence even in dynamic scenes. The eclipse footage example reveals VideoPASTA’s ability to describe gradual visual transformations accurately, avoiding the baseline’s tendency to oversimplify temporal transitions. Finally, the instruction scene identifying magic demonstrates our model’s capability to establish clear causal relationships between actions and their outcomes, supported by specific visual evidence.
@@ Appendix G Prompt Templates
Preference Data Filtering Prompt You have a single video input and a set of four responses for each query: 1. One aligned response that is claimed to be well-aligned with the video content. 2. Three adversarial responses, each intentionally introducing spatial, temporal, or cross-frame errors. The goal is to validate that: The aligned response truly aligns with the query (no unintended contradictions or inaccuracies). Each adversarial response introduces a clear misalignment without merely restating or slightly rephrasing the aligned response. For each query and its four responses: 1. Sanity-check the aligned response. Confirm that it accurately reflects the video’s content in relation to the query. If any errors or contradictions are detected, discard them. 2. Examine each adversarial response. Identify whether it deliberately contradicts or distorts the query/video content (e.g., reversed sequence, false spatial claims). If it is too similar to the aligned response or fails to demonstrate a clear misalignment, discard it.
Figure 12: Prompt template for validating one aligned and three adversarial responses to ensure robust preference pairs.
@@
Adversarial Example Evaluation Prompt Task: Evaluate whether the provided adversarial example correctly targets its intended failure mode in video understanding. Query: [Original question asked about the video] Aligned Response: [The correct/preferred response to the query] Adversarial Example: [The adversarial example to be evaluated] Claimed Failure Mode: [One of: “Spatial Misalignment”, “Temporal Incoherence”, or “Cross-Frame Disconnection”] Failure Mode Definitions: Spatial Misalignment: Incorrectly describing spatial relations, object positions, occlusion patterns, depth, or relative positioning within a single frame. Temporal Incoherence: Violating the natural ordering of events, describing sequential actions as simultaneous, merging distinct events, or misordering the sequence of activities shown in the video. Cross-Frame Disconnection: Breaking object persistence across frames, describing the same object as different entities across scenes, failing to maintain character/object consistency, or incorrectly describing changes between distant frames. Evaluation Instructions: 1. Carefully analyze the adversarial example in relation to the aligned response. 2. Determine if the adversarial example genuinely induces the claimed failure mode. 3. Your evaluation should be based solely on the definitions provided above. 4. Provide a binary judgment: “Yes” if the adversarial example correctly targets the claimed failure mode, “No” if it does not. 5. Briefly explain your reasoning (2-3 sentences). Output Format: